OpenStudy (anonymous):

An aqueous solution freezes at -3.47⁰C. What is its boiling point?

OpenStudy (anonymous):

@JFraser @Abmon98 @kavonta @bibby @micahm @Tj1014

OpenStudy (anonymous):

101.13°C

OpenStudy (anonymous):

how so? i was told this: The molar freezing point depression constant for water is 1.86°C/mol. You have a solution that freezes at -3.47°C. Therefore it is 3.47 /1.86 = 1.9 molar solution The boiling point elevation for water is 0.512°C/mol You have a 1.9m solution, so the solution will have boiling point elevation = 0.512*1.9 = °C

OpenStudy (anonymous):

The molal freezing point depression constant for water is 1.86°C/mol. You have a solution that freezes at -4.09°C. Therefore it is 4.09/1.86 = 2.2 molal solution The boiling point elevation for water is 0.512°C/mol You have a 2.2m solution, so the solution will have boiling point elevation = 0.512*2.2 = 1.13°C

OpenStudy (micahm):

he right

OpenStudy (anonymous):

i dont have that question tho. its not -4.09 degrees its -3.47 degrees

OpenStudy (micahm):

so what are you looking for

OpenStudy (anonymous):

the boiling point of -3.47 Degrees

OpenStudy (micahm):

On the Fahrenheit scale, the freezing point of water is 32 degrees Fahrenheit (°F) and the boiling point is 212 °F (at standard atmospheric pressure). This puts the boiling and freezing points of water exactly 180 degrees apart.

OpenStudy (anonymous):

alright.

OpenStudy (micahm):

do you need more info

OpenStudy (anonymous):

i need more help to solve the problem : 3.47/1.86=1.86

OpenStudy (anonymous):

then 0.512*1.86 = 0.95232°C ?

OpenStudy (anonymous):

@JFraser can you please help me?

OpenStudy (micahm):

is this a math problem if so i have a site that can help you

OpenStudy (micahm):

www.freemathhelp.com

OpenStudy (anonymous):

its a chemistry question.

OpenStudy (jfraser):

use the freezing point depression to find the molality of the solution, as @Tj1014 says. then plug that same molality into the boiling point expression. the only difference between the two formulas is the value of K. The Kf of water is -1.86C/molal, but the Kb of water is 0.512C/molal \[\Delta T_f = K_f*m\]and\[\Delta T_b = K_b*m\]

OpenStudy (anonymous):

delta TF = -1.86*m delta TB= 0.512*m What is M?

OpenStudy (jfraser):

No. \(\Delta T_f\) is -3.47C. The problem tells you that. Use the first set of information to find m, then plug that m into the second equation and solve for \(\Delta T_b\)

OpenStudy (anonymous):

so -3.47*-1.86 to find delta TF

OpenStudy (jfraser):

no. it tells you \(\Delta T_f\) = -3.47C. Solve the first equation for \(m\)